The pedals of a bicycle rotate in a circle with a diameter of 50cm. What is the maximum torque a 75kg rider can apply by putting all of her weight on one pedal?
1
Expert's answer
2011-11-17T12:05:29-0500
The rider applies a force F = mg when she puts all the weight on the
pedal.
Using the formula: torque = r*F, where r is the distance from the axis
(radius = d/2), we get:
torque = m*g*d/2 = 75*9.8*0.5/2 = 183.75 (H*m).